The coconut demand function is Q=1200-9.5p+16.2pp+0.2y . Assume that p is initially $ 0.45 per kg, pp =$ 0.31 per kg and Q=1275 thousand matric tons per year. Calculate income elasticity of demand for coconut oil?
1
Expert's answer
2015-06-09T11:20:49-0400
Qd = 1200-9.5p+16.2pp+0.2y.
If p is initially $ 0.45 per kg, pp =$ 0.31 per kg and Q=1275 thousand
metric tons per year, then 1275 = 1200 - 4.275 + 5.022 + 0.2y, 
0.2y = 74.253
y = 371.265
Income elasticity of demand for coconut oil is Qd'(y) = (1200 - 4.275 + 5.022 + 0.2y)' = 0.2.